# publications.bib

### v3

@@ Appendix A Spectral Decomposition of Sheaves on $\mathrm{Bun}_{G}$, by David Hansen
Let $G/\mathbb{Q}_{p}$ be a connected reductive group, $\Lambda/\mathbf{Z}_{\ell}$ an algebraically closed field. If $\mathrm{char}(\Lambda)\neq 0$ we assume $\ell$ is very good for $G$.
Set $\mathrm{D}(\mathrm{Bun}_{G})=\mathrm{D}_{\mathrm{lis}}(\mathrm{Bun}_{G},\Lambda)$ to be the derived category of lisse-étale $\Lambda$-sheaves, regarded as a stable $\infty$-category whenever convenient. Let $\mathfrak{X}_{\hat{G}}=Z^{1}(W_{E},\hat{G})_{\Lambda}/\hat{G}$ be the stack of $L$-parameters over $\Lambda$, and let $X_{\hat{G}}$ be its coarse moduli space, $q:\mathfrak{X}_{\hat{G}}\to X_{\hat{G}}$ the natural map. We will regard $\mathfrak{X}_{\hat{G}}$ as a disjoint union of finite type algebraic stacks over $\Lambda$, and $X_{\hat{G}}$ as a disjoint union of finite type affine $\Lambda$-schemes. As in [FS], we have the spectral action of $\mathrm{Perf}(\mathfrak{X}_{\hat{G}})$ on $\mathrm{D}(\mathrm{Bun}_{G})$, and there is a natural map $\Psi_{G}:\mathcal{O}(\mathfrak{X}_{\hat{G}})=\mathcal{O}(X_{\hat{G}})\to\mathfrak{Z}(\mathrm{D}(\mathrm{Bun}_{G})):=\pi_{0}(\mathrm{id}_{\mathrm{D}(\mathrm{Bun}_{G})})$, where we recall that $Z^{1}(W_{E},\hat{G})_{\Lambda}$ is a disjoint union of affine schemes by [FS, Theorem VIII.1.3]. These two structures are compatible (as proven by Zou [Zou, Theorem 5.2.1]).
By [FS, Prop. VIII.3.8], the set of closed points $X_{\hat{G}}(\Lambda)$ is naturally in bijection with the set of isomorphism classes of semisimple $L$-parameters $\phi:W_{E}\to\phantom{}^{L}\!G(\Lambda)$. Let $\mathfrak{m}_{\phi}\subset\mathcal{O}(X_{\hat{G}})$ be the maximal ideal associated with a given $\phi$.
@@
By construction, $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ is a full subcategory of $\mathrm{D}(\mathrm{Bun}_{G})$ stable under arbitrary limits and colimits, and the tautological inclusion functor $\iota_{\phi}:\mathrm{D}(\mathrm{Bun}_{G})_{\phi}\hookrightarrow\mathrm{D}(\mathrm{Bun}_{G})$ commutes with limits and colimits. By the $\infty$-categorical adjoint functor theorem [HTT, Cor. 5.5.2.9.(2)], it therefore admits a left adjoint $\mathcal{L}_{\phi}:\mathrm{D}(\mathrm{Bun}_{G})\to\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$. [^10] To see that $\iota_{\phi}$ is accessible, use [HTT, Prop. 5.4.7.7] together with the fact that $\iota_{\phi}$ admits a right adjoint, which follows from [HTT, Cor. 5.5.2.9.(1)]. The unit of the adjunction gives a map $A\to\iota_{\phi}\mathcal{L}_{\phi}A=:A_{\phi}$ functorially in $A$. Since $\iota_{\phi}$ is fully faithful, $\mathcal{L}_{\phi}\iota_{\phi}=\mathrm{id}$, so $(A_{\phi})_{\phi}=A_{\phi}$, i.e. the endofunctor $A\rightsquigarrow A_{\phi}$ is idempotent. We remark that $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ is a Bousfield localization of $\mathrm{D}(\mathrm{Bun}_{G})$, and the map $A\to A_{\phi}$ is the initial map from $A$ to a $\phi$-local sheaf.
@@
The full subcategory $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ is preserved by the spectral action, and $A\rightsquigarrow A_{\phi}$ commutes with the spectral action. Moreover, $\mathrm{supp}(A_{\phi})\subseteq\mathrm{supp}(A)$.
@@
The first claim is clear, since the spectral action commutes with the action of $\mathcal{O}(X_{\hat{G}})$. For the remaining claims (and some later arguments), it is useful to give an explicit formula for $A_{\phi}$. Let $\mathcal{I}_{\phi}$ be the diagram category whose objects are elements of $\mathcal{O}(X_{\hat{G}})\smallsetminus\mathfrak{m}_{\phi}$ and where a morphism $f\to g$ is an element $h\in\mathcal{O}(X_{\hat{G}})\smallsetminus\mathfrak{m}_{\phi}$ such that $g=fh$. This is clearly cofiltered. Let $F\in\mathrm{Fun}(\mathcal{I}_{\phi},\mathrm{D}(\mathrm{Bun}_{G}))$ be the functor sending $f$ to $A$ and sending a morphism $h\in\mathrm{Mor}(f,g)$ to $\cdot h\in\mathrm{End}(A)$. Then $A_{\phi}=\mathrm{colim}_{i\in\mathcal{I}_{\phi}}F(i)$. The remaining claims are now immediate. ∎
To make sense of the next proposition, note that for any $A,B\in\mathrm{D}(\mathrm{Bun}_{G})$, $\mathrm{Hom}(B,A)$ is naturally a $\mathfrak{Z}(\mathrm{D}(\mathrm{Bun}_{G}))$-module, whence a $\mathcal{O}(X_{\hat{G}})$-module.
@@
If $C\in\mathrm{D}(\mathrm{Bun}_{G})$ is compact, then $\mathrm{Hom}(C,A_{\phi})\cong\mathrm{Hom}(C,A)_{\mathfrak{m}_{\phi}}$ functorially in $A$ and $C$, where the RHS is the usual localization as an $\mathcal{O}(X_{\hat{G}})$-module.
@@
Recall from [FS, Prop. VII.7.9] that $B\in\mathrm{D}(\mathrm{Bun}_{G})$ is ULA iff $R\mathrm{Hom}(C,B)\in\mathrm{Perf}(\Lambda)$ is a perfect complex for all compact objects $C\in\mathrm{D}(\mathrm{Bun}_{G})$. Now, if $C$ is compact, $R\mathrm{Hom}(C,-)$ commutes with filtered colimits, so
@@
is an isomorphism, since $\mathrm{D}(\mathrm{Bun}_{G})$ is compactly generated [FS, Theorem I.5.1 (iii)]. As in the previous proof, $R\mathrm{Hom}(C,A)$ is a perfect complex, so $\mathrm{Hom}(C,A)$ is a finite $\Lambda$-vector space. In particular, it is a finite length $\mathcal{O}(X_{\hat{G}})$-module supported at a finite set of closed points $S\subset X_{\hat{G}}(\Lambda)$, so if $\phi\notin S$ then $\mathrm{Hom}(C,A_{\phi})=\mathrm{Hom}(C,A)_{\mathfrak{m}_{\phi}}=0$ using Proposition A.3. We then conclude that
@@
commutes, where $f\mapsto f^{\vee}$ is the involution of $\mathcal{O}(X_{\hat{G}})$ induced by composition with the Chevalley involution at the level of $L$-parameters. Since $f\in\mathfrak{m}_{\phi}$ iff $f^{\vee}\in\mathfrak{m}_{\phi^{\vee}},$ we deduce that if $A$ is $\phi$-local then $\mathbb{D}_{\mathrm{Bun}_{G}}(A)$ is $\phi^{\vee}$-local. Using biduality, we also get that if $A$ is ULA then $A$ is $\phi$-local if and only if $\mathbb{D}_{\mathrm{Bun}_{G}}(A)$ is $\phi^{\vee}$-local.

### v5

@@ Appendix A Spectral Decomposition of Sheaves on $\mathrm{Bun}_{G}$, by David Hansen
Let $G/\mathbb{Q}_{p}$ be a connected reductive group, $\Lambda/\mathbf{Z}_{\ell}$ an algebraically closed field. If $\mathrm{char}(\Lambda)\neq 0$ we assume $\ell\nmid|\pi_{0}(Z(G))|$, where $Z(G)$ denotes the center of $G$, as in [16, Theorem I.10.1].
[…]
@@
By construction, $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ is a full subcategory of $\mathrm{D}(\mathrm{Bun}_{G})$ stable under arbitrary limits and colimits, and the tautological inclusion functor $\iota_{\phi}:\mathrm{D}(\mathrm{Bun}_{G})_{\phi}\hookrightarrow\mathrm{D}(\mathrm{Bun}_{G})$ commutes with limits and colimits. Since $\mathrm{D}(\mathrm{Bun}_{G})$ is presentable, $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ is then presentable by [51, Theorem 1.1]. By the $\infty$-categorical adjoint functor Theorem [45, Cor. 5.5.2.9.(2)], the inclusion $\iota_{\phi}$ therefore admits a left adjoint $\mathcal{L}_{\phi}:\mathrm{D}(\mathrm{Bun}_{G})\to\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$. [^15] To see that $\iota_{\phi}$ is accessible, use [45, Prop. 5.4.7.7] together with the fact that $\iota_{\phi}$ admits a right adjoint, which follows from [45, Cor. 5.5.2.9.(1)]. The unit of the adjunction gives a map $A\to\iota_{\phi}\mathcal{L}_{\phi}A=:A_{\phi}$ functorially in $A$. Since $\iota_{\phi}$ is fully faithful, $\mathcal{L}_{\phi}\iota_{\phi}=\mathrm{id}$, so $(A_{\phi})_{\phi}=A_{\phi}$, i.e. the endofunctor $A\rightsquigarrow A_{\phi}$ is idempotent. We remark that $\mathrm{D}(\mathrm{Bun}_{G})_{\phi}$ is a localization of $\mathrm{D}(\mathrm{Bun}_{G})$, and the map $A\to A_{\phi}$ is the initial map from $A$ to a $\phi$-local sheaf.
@@
[…]
Here $\mathrm{supp}(A)\subset|\mathrm{Bun}_{G}|=B(G)$ denotes the set of points $b$ such that $i_{b}^{\ast}A\neq 0$.
@@
[…]
@@
[…]
@@
Recall from [16, Prop. VII.7.9] that $B\in\mathrm{D}(\mathrm{Bun}_{G})$ is ULA iff $R\mathrm{Hom}(C,B)\in\mathrm{Perf}(\Lambda)$ is a perfect complex for all compact objects $C\in\mathrm{D}(\mathrm{Bun}_{G})$. Now, if $C$ is compact, $R\mathrm{Hom}(C,-)$ commutes with filtered colimits, so
@@
[…]
@@
[…]